Documenta Mathematica

Journal der Deutschen Mathematiker-Vereinigung

Band 2, 1997

A. Böttcher On the Approximation Numbers

of Large Toeplitz Matrices  1–29

Amnon Besser On the Finiteness of � for Motives

Associated to Modular Forms  31–46

A. Langer Selmer Groups and Torsion Zero Cycles on the

Selfproduct of a Semistable Elliptic Curve  47–59

Christian Leis Hopf-Bifurcation in Systems with Spherical Symmetry

Part I : Invariant Tori  61–113

Jane Arledge, Marcelo Laca and Iain Raeburn Semigroup Crossed Products and Hecke Algebras

Arising from Number Fields  115–138

Joachim Cuntz Bivariante K-Theorie für lokalkonvexe Algebren

und der Chern-Connes-Charakter  139–182

Henrik Kratz Compact Complex Manifolds

with Numerically Effective Cotangent Bundles  183–193

Ekaterina Amerik

Maps onto Certain Fano Threefolds  195–211

Jonathan Arazy and Harald Upmeier Invariant Inner Product in Spaces of Holomorphic Functions

on Bounded Symmetric Domains  213–261

Victor Nistor Higher Index Theorems and

the Boundary Map in Cyclic Cohomology  263–295

Oleg T. Izhboldin and Nikita A. Karpenko

On the Group H^{3}(F (ψ, D)/F)  297–311

Udo Hertrich-Jeromin and Franz Pedit Remarks on the Darboux Transform of

Isothermic Surfaces  313–333

Udo Hertrich-Jeromin Supplement on Curved Flats in the Space of Point Pairs and Isothermic Surfaces:

A Quaternionic Calculus  335–350

Ernst-Ulrich Gekeler On the Cuspidal Divisor Class Group of a

Drinfeld Modular Curve  351–374

Mikael Rørdam

Stability of C^{∗}-Algebras is Not a Stable Property  375–386